OpenStudy (anonymous):

A function is given below. Determine the average rate of change of the function between x = -3 and x = -3 + h. g(x)=-10/x

OpenStudy (anonymous):

average rate of change {g(-3+h)- g(-3)}/ h

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

i know the formula but some how the answer that i get is wrong... I just need help working it out

OpenStudy (anonymous):

so let's do the calculation here or post ur work here.

OpenStudy (anonymous):

i got \[(-10/(-3+h))-(5/3)/(h)\]

OpenStudy (anonymous):

g(-3)=10/3

OpenStudy (anonymous):

g(-3+h)= -10/(-3+h)

OpenStudy (anonymous):

i see that the g(-3) is one thing i got wrong

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then after i calculated it again i got \[ \frac{ -h }{ -9+3h }\]

OpenStudy (anonymous):

which is still wrong :/

OpenStudy (anonymous):

{-10/(-3+h) -10/3} /h =-10/h { 1/(-3+h) +1/3 } =-10/h [ h/ {(-3+h) 3}] = -10/{(-3+h)3}

OpenStudy (anonymous):

i also got that and it's saying its wrong?

OpenStudy (anonymous):

u got -h/ (-9+3h) but I got 10/(9-3h)

OpenStudy (anonymous):

what's the correct answer?

OpenStudy (anonymous):

i kept putting in extra negatives...

OpenStudy (anonymous):

and btw i am doing my homework through a website that auto checks my work and i have 5 tries to get it right